OpenStudy (anonymous):

IF A is (0,0) and C is (a,a) what is the slope of AC?

jimthompson5910 (jim_thompson5910):

slope: m = (y2-y1)/(x2-x1) m = (a-0)/(a-0) m = a/a m = 1 So the slope of AC is 1.

OpenStudy (anonymous):

slope would be 1